OpenStudy (anonymous):

TEST TOMORROW!! help is much appreciated!!! Linear Algebra: (2x-1) (5x+1/2)=0 and this problem: y=x^3-x xe{0,1,2,3,4} <--graph

OpenStudy (anonymous):

helping with any one problem will help a ton :[

myininaya (myininaya):

linear algebra or college algebra? (2x-1)(5x+1/2)=0 2x-1=0 or 5x+1/2=0 x=1/2 or x=-1/10

myininaya (myininaya):

i don't understand what you are asking in second problem

myininaya (myininaya):

\[f(x)=x^3-x \] \[f(0)=0^3-0=0; f(1)=1^3-1=0;f(2)=2^3-2=6;f(3)=3^3-3=24\]

myininaya (myininaya):

\[f(4)=4^3-4-60\]

OpenStudy (anonymous):

xe stands for x intercept i think. its a greek symbol for that...and then i think i have to subsitute with those numbers...not exactly sure. and thank you

myininaya (myininaya):

\[(0,0);(1,0);(2,6);(3,24)\]

myininaya (myininaya):

i guess you just wanted the coordinates? thats what i was unsure about

OpenStudy (anonymous):

yeah it said to graph it so im assuming thats correct. thankyou! :)